OpenStudy (anonymous):

Factor completely 3ab − 5ax + 9b − 15x.

OpenStudy (anonymous):

@iyuko

OpenStudy (anonymous):

srry the server is going so slow

iYuko (iyuko):

3ab-5ax+9b-15x = 3ab+9b-5ax-15x = 3b(a+3)-5x(a+3) = (3b-5x)(a+3)

iYuko (iyuko):

I can help with more just tag me

OpenStudy (anonymous):

ill tag u in the next one

iYuko (iyuko):

k